# Imaging of I Zw 18 by JWST: I. Strategy and First Results of Dusty Stellar Populations

### v1

@@ 2.1 NIRCam Strategy
Our NIRCam observations were taken with the full frame and without need for a primary dither pattern, as no mosaic gaps needed to be filled in. The shallow4 readout pattern was selected in effort to optimize the signal-to-noise ratio (S/N), with one integration per exposure and six groups per integration. In addition, we implemented a four-point sub-pixel dither pattern to sample the point spread function (PSF), which resulted in an overall exposure time of 1245.5 seconds for each of the SW + LW filter combinations (i.e., F115W + F356W and F200W + F444W).
Our selection of NIRCam filters was based on prior work on stellar populations at these wavelengths. The F115W and F200W filters are similar to standard Johnson J and K, respectively, while the F356W and F444W filters are similar to Spitzer IRAC [3.6] and [4.5], respectively. This allows for comparative studies with previous observations made of other galaxies, including the SAGE studies of the Magellanic Clouds (e.g., Blum et al., 2006; Meixner et al., 2006; Whitney et al., 2008; Gordon et al., 2011; Boyer et al., 2011), the DUSTiNGS project which observed many nearby SF galaxies (e.g., Boyer et al., 2015a, b; McQuinn et al., 2017; Goldman et al., 2019a), and others (e.g., Nikolaev & Weinberg, 2000; Jones et al., 2015a, 2018; Goldman et al., 2019b; Jones et al., 2019; Hirschauer et al., 2020).
Centering the spatial extent of I Zw 18 within the upper-left quadrant of Module B (B3) was accomplished by requesting a special requirement which offset the position of the target by 55.0 arcsec in the “X” direction and 35.0 arcsec in the “Y” direction, centered at RA = 09:34:02.10, Dec. = +55:14:25.00. This allows for uninterrupted spatial coverage of I Zw 18 in the NIRCam SW FoV without relying on a restriction in spacecraft position angle.
@@ 2.2 MIRI Strategy
Observations with MIRI employed a four-point cycling dither pattern and a small pattern size, utilizing the full subarray and fastr1 readout pattern, centered at RA = 09:34:02.00, Dec. = +55:14:28.00. For the F770W and 1000W filters, we implemented strategies of 100 and 200 groups per integration, respectively, with one integration per exposure, for four total integrations each. The total exposure time for the F770W filter was 1110.0 seconds, while for F1000W, it was 2220.0 seconds. With the next-longest wavelength filter, F1500W used 110 groups per integration and two integrations per exposure, resulting in eight total integrations totaling 2453.1 seconds. The longest-wavelength MIRI filter, F1800W, necessitated selecting 88 groups per integration to avoid saturation of the sky background, achieving a long total exposure time of 5916.4 seconds by increasing the number of integrations per exposure to six, for 24 total integrations.
Our selection of MIRI filters was guided by the predicted JWST fluxes and color separations for IR-bright populations from Jones et al. (2017). Color-magnitude diagrams (CMDs) utilizing these long wavelengths are sensitive to the reddest and dustiest sources, offering unambiguous identification of evolved (RSGs and AGB stars) and young (YSOs) stellar populations via long wavelength baselines. The longer-wavelength MIRI filters F2100W and F2550W were not chosen despite the wider possible baselines of its CMDs, owing to a steep drop-off in sensitivity which would have required substantially longer integration times to achieve comparable S/N. Separation of source types was determined to be equivalently achievable using the F1800W filter with shorter exposure times. Finally, detection of polycyclic aromatic hydrocarbon (PAH) emission is accomplished with the F770W filter, and the F1000W filter is sensitive to the 10 m silicate feature.
In addition, a MIRI coordinated parallel observation (when NIRCam was prime) was included to provide background comparison images useful for contaminant source subtraction. The MIRI coordinated parallel image was obtained in the F1000W and F1500W filters utilizing a full subarray, and designed to maximize available integration time as determined by the NIRCam prime observations. We utilized a fastr1 readout pattern, with 112 groups per integration and one integration per exposure over the four total dithers, resulting in a total exposure time of 1243.2 seconds per filter. Data acquired from the MIRI parallel field is not considered further in this paper.
@@ 3 Data Reduction and Source Extraction
3.1 Data Reduction Methodology
Processing of the NIRCam imaging data was carried out utilizing a slightly modified version of the JWST official pipeline (version 1.7.2) through all three stages. These modifications corrected for 1/f noise and flat field correction noise, World Coordinate System (WCS) alignment issues, and included the most recent NIRCam calibration files at the time of data processing. For Stage-1 data, we used jwst_1069.pmap of the Operational Pipeline Calibration Reference Data System (CRDS), produced on 3 October 2022 with on-sky derived photometric zero-points (Gordon et al., 2022; Boyer et al., 2022). A frame0 correction was implemented to recover stars saturated in the first group (21 seconds), but were unsaturated in the first 10.7 seconds comprising frame0 (ramp_fit.suppress_one_group=False). For Stage-2 output files (*_cal.fits), image1overf.py111https://github.com/chriswillott/jwst (Willott, 2022) was used to correct for 1/f noise. Because the small field of view contained few suitable Gaia reference stars, astrometric alignment was done relatively between frames and filters. A simple source catalog was created from a single arbitrary exposure in the F115W filter using source detection from the JWST pipeline. Alignment of the remaining F115W frames and frames from the remaining filters to this source catalog was accomplished using the JWST/Hubble Alignment Tool (JHAT222https://github.com/arminrest/jhat; Rest, 2023). Because alignment to Gaia was already accomplished via JHAT, the Stage-3 tweakreg step was skipped.
@@ 3.1 Data Reduction Methodology
JWST pipeline version 1.12.0 with CRDS version 11.17.9 and context jwst_1149.pmap were used in the creation of the final F770W imaging data. The remaining MIRI filters were subsequently processed using CRDS version 11.17.6 and context jwst_1179.pmap. The new CRDS context included updates for the MIRI Medium Resolution Spectrometer and thus does not affect the MIRI imaging data we present here. Flux-calibrated images in all four filters were produced by processing the raw MIRI ramp files using Dectector1Pipeline, with the output being run through Image2Pipeline with default parameters. Before combining the individual dithers into a final image, we aligned the F770W Image2Pipeline files (*_cal.fits) to the NIRCam F444W point-source catalog. We then applied the same astrometric corrections to the remaining MIRI filters. Following image alignment, we perform instrumental background subtraction following the method outlined in Nally et al. (2023). For each filter, we first median combine all dithers. Because I Zw 18 is present in each of these exposures, however, the median combined backgrounds are not free of structure. We therefore select unaffected detector rows and columns and map these to the detector plane to produce a model background, which we then subtract from each individual exposure (see Nally et al., 2023, Dicken et al. submitted). Finally, the background-subtracted and F444W-aligned individual dither exposures were combined into final images using Image3Pipeline with the tweakreg step turned off.
3.2 Point Source Extraction
@@
The extraction of point sources in I Zw 18 was accomplished through use of the starbugii (Nally, 2023) photometric tool and pipeline. This software is optimized for JWST observations of crowded stellar fields within complex environments, performing point-source extraction and band merging across multiple observations and wavelengths, utilizing core functions from the python photutils (Bradley et al., 2022) package. Examples of programs which have employed starbugii for photometric study include Jones et al. (2023), Lenkić et al. (2023), Nally et al. (2023), and Habel et al. (submitted). The complete set of the relevant starbugii parameters and their adopted values used for this program’s photometric extractions are listed in Table 3.
@@
For each filter in NIRCam, we perform source detection using starbug2 detect on the Stage-3 drizzled image. We locate sources that are above the sky level and their location is measured by fitting centroids. Values of the sharpness and roundness are assigned en route to calculating a geometry for each point source. By setting an upper limit on sharpness, the measured ratio of source peak height to its median pixel value, remaining image artefacts are removed from the catalog, while setting a lower limit assists in removing faint peaks in the dust structure. Source symmetry is measured by the roundness parameter. Most resolved background galaxies, as well as further spurious detections within the dust structure, are removed by limiting the allowed level of asymmetry in the source profiles. The same routine was used to detect sources in the F770W MIRI image with modified detection parameters. Because the decrease in image resolution inhibited the detection of robust point-like sources in the longer-wavelength filters (F1000W, F1500W, and F1800W), bright sources in close proximity to other similar objects were instead merged into bright clustered regions that were not analyzed in this study.
@@
Aperture photometry was performed for all four NIRCam filters (F115W, F200W, F356W, and F444W), plus the shortest-wavelength MIRI filter (F770W). For all NIRCam sources, we use a fixed radius of 1.5 pixels as well as a sky annulus of 3.0 and 4.5 pixels for the inner and outer radii, respectively. Aperture correction is interpolated between values given in CRDS jwst_nircam_apcorr_0004.fits. For the MIRI F770W band, the procedure is the same, however the aperture radius is increased to two pixels to account for the larger PSFs intrinsic to longer wavelengths, backgrounds are calculated within annuli of four and six pixels for the inner and outer radii, respectively, and jwst_miri_apcorr_0005.fits is used to calculate aperture corrections. Through inspection of the data quality array within the aperture of each source, we flag those with saturated, DO_NOT_USE, or SRC_JMP pixels (indicating a jump during detection) with the SRC_BAD flag in starbugii.
@@
We perform PSF photometry on the stellar sources of I Zw 18 on the individual Stage-2 exposures of all four NIRCam filters. First, the nebulous emission which underlies our imaging data is modelled with the routine starbug2 background. This places masking apertures of variable size, and fills them with the median pixel value within a local annulus. Values of the diffuse emission background are then measured for every pixel in the image by averaging all local pixels within a set box size, creating an effective representation of the nebulous emission. A clean image of the field, with nebulous emission estimate subtracted, is therefore created which enables accurate fitting of PSFs.
The tool webbpsf (Perrin et al., 2014) version 1.1.1 was then employed to generate a five-arcsecond PSF for each NIRCam detector subarray. After estimating and subtracting the nebulous background from each single exposure, PSF photometry was performed by running starbug2 psf on the background-subtracted image, at locations defining source positions determined via the source detection. By leaving the centroid position as a free parameter, both the flux and position of each source are allowed to be fit. If the position of the new centroid differs from the initial guess determined during the source detection step by more than 0.1 arcseconds, the flux is re-fit with the source’s position held fixed to the initial guess and is flagged with SRC_FIX, denoting the fit as being of poorer quality.
The calculation and application of instrumental zero point magnitudes are necessary to calibrate the PSF photometry, as they are not normalized to physical units. A median difference in source magnitudes measured by the aperture photometry and PSF fitting is calculated using starbug2 calc-instr-zp, allowing for the determination of instrumental zero point magnitudes, which are then employed to calibrate the PSF photometry from starbugii to the AB magnitude system. These magnitudes are then converted to the Vega system using the reference file jwst_nircam_abvegaoffset_0002.asdf.
@@ 3.3 Stellar Source Catalog
Point-source catalogs for each of the four NIRCam filters were created via the methods described in §3.2. A band-matched catalog of point sources is then constructed for I Zw 18 by merging these catalogs produced for the individual filters using starbug2-match band. The possibility of mismatching high-quality stellar data with any spurious detector artefacts is mitigated by excluding all sources in the individual filter catalogs with an error on their magnitude greater than 10%: While this may limit the potential for completeness of the master catalog, this strategy allows for better reliability overall.
First, we consider the catalog made from the shortest-wavelength filter data, for which the smaller PSF full width at half maximum (FWHM) possesses astrometry of the highest certainty. These sources are then matched using a nearest neighbor algorithm to sources in the next-shortest wavelength catalog. The threshold for separation between filter catalogs increases with increasing wavelength (equivalently, as the size of the PSF increases), as adopting a single separation threshold which is larger than the astrometric uncertainty of, for example, the longest-wavelength NIRCam filter, would result in mismatches with, for example, the shortest-wavelength NIRCam filter. For NIRCam, the SW filters (F115W and F200W) are matched using a threshold of 0.06″, while the LW filters (F356W and F444W) are matched and added using a threshold of 0.10″. Any unmatched sources outside of the separation threshold are then appended to the end of the catalog, before the process is repeated with the next-shortest wavelength filter data which remains. We note that, for the nearest neighbor search method, the position coordinates for the source are adopted from the shortest-wavelength catalog from which said source initially appears. This allows us to retain the greatest possible positional accuracy for each source.
@@
To avoid artifacts at the edges of the detectors and unresolved background galaxies impacting our photometric analysis, an elliptical region approximately 55″40″ aligned along I Zw 18’s semi-major axis was defined. Only point sources detected within this region ( = 1,520; see gray points from Figure 6) were considered to be members of I Zw 18 in our analysis. In addition, point sources from Component C were extracted from within a 30″22″ elliptical region centered on this system, while carefully avoiding the detector edge. These points are likewise shown in Figure 6 as red open circles ( = 286). The proximity of Component C to the detector edge makes robust separation between real sources and artifacts challenging, here we present preliminary results in §4.3, with a more detailed analysis to follow as the JWST calibrations improve.

### v2

@@ 2.1 NIRCam Strategy
Our NIRCam observations were taken with the full frame and without need for a primary dither pattern, as there were no mosaic gaps that needed to be filled in. The shallow4 read-out pattern was selected in effort to optimize the signal-to-noise ratio (S/N), with one integration per exposure and six groups per integration. In addition, we implemented a four-point subpixel dither pattern to sample the point spread function (PSF), which resulted in an overall exposure time of 1245.5 s for each of the SW+LW filter combinations (i.e., F115W+F356W and F200W+F444W).
Our selection of NIRCam filters was based on prior work on stellar populations at these wavelengths. The F115W and F200W filters are similar to standard Johnson J and Ks, respectively, while the F356W and F444W filters are similar to Spitzer IRAC [3.6] and [4.5], respectively. This allows for comparative studies with previous observations made of other galaxies, including the SAGE studies of the Magellanic Clouds (e.g., Blum et al., 2006; Meixner et al., 2006; Whitney et al., 2008; Gordon et al., 2011; Boyer et al., 2011), the DUSTiNGS project, which observed many nearby SF galaxies (e.g., Boyer et al., 2015a, b; McQuinn et al., 2017; Goldman et al., 2019a), and others (e.g., Nikolaev & Weinberg, 2000; Jones et al., 2015a, 2018; Goldman et al., 2019b; Jones et al., 2019; Hirschauer et al., 2020).
Centering the spatial extent of I Zw 18 within the upper-left quadrant of Module B (B3) was accomplished by requesting a special requirement which offset the position of the target by 55″.0 in the “X” direction and 35″.0 in the “Y” direction, centered at R.A. = 09:34:02.10, decl. = +55:14:25.00. This allows for uninterrupted spatial coverage of I Zw 18 in the NIRCam SW FOV without relying on a restriction in spacecraft position angle.
@@ 2.2 MIRI Strategy
Observations with MIRI employed a four-point cycling dither pattern and a small pattern size, utilizing the full subarray and fastr1 read-out pattern, centered at R.A. = 09:34:02.00, decl. = +55:14:28.00. For the F770W and 1000W filters, we implemented strategies of 100 and 200 groups per integration, respectively, with one integration per exposure, for four total integrations each. The total exposure time for the F770W filter was 1110.0 s, while for F1000W it was 2220.0 s. With the next-longest-wavelength filter, F1500W used 110 groups per integration and two integrations per exposure, resulting in eight total integrations totaling 2453.1 s. The longest-wavelength MIRI filter, F1800W, necessitated selecting 88 groups per integration to avoid saturation of the sky background, achieving a long total exposure time of 5916.4 s by increasing the number of integrations per exposure to six, for 24 total integrations.
Our selection of MIRI filters was guided by the predicted JWST fluxes and color separations for IR-bright populations from Jones et al. (2017). Color–magnitude diagrams (CMDs) utilizing these long wavelengths are sensitive to the reddest and dustiest sources, offering unambiguous identification of evolved (RSGs and AGB stars) and young (YSOs) stellar populations via long-wavelength baselines. The longer-wavelength MIRI filters F2100W and F2550W were not chosen despite the wider possible baselines of its CMDs, owing to a steep drop-off in sensitivity which would have required substantially longer integration times to achieve comparable S/N. Separation of source types was determined to be equivalently achievable using the F1800W filter with shorter exposure times. Finally, detection of polycyclic aromatic hydrocarbon (PAH) emission is accomplished with the F770W filter, and the F1000W filter is sensitive to the 10 m silicate feature.
In addition, a MIRI coordinated parallel observation (when NIRCam was prime) was included to provide background comparison images useful for contaminant source subtraction. The MIRI coordinated parallel image was obtained in the F1000W and F1500W filters utilizing a full subarray, and designed to maximize available integration time as determined by the NIRCam prime observations. We utilized a fastr1 read-out pattern, with 112 groups per integration and one integration per exposure over the four total dithers, resulting in a total exposure time of 1243.2 s per filter. Data acquired from the MIRI parallel field are not considered further in this paper.
@@ 3 Data Reduction and Source Extraction
3.1 Data-reduction Methodology
Processing of the NIRCam imaging data was carried out utilizing a slightly modified version of the JWST official pipeline (version 1.7.2) through all three stages. These modifications corrected for 1/f noise and flat-field correction noise, World Coordinate System (WCS) alignment issues, and included the most recent NIRCam calibration files at the time of data processing. For Stage-1 data, we used jwst_1069.pmap of the Operational Pipeline Calibration Reference Data System (CRDS), produced on 2022 October 3 with on-sky derived photometric zero-points (Boyer et al., 2022; Gordon et al., 2022). A frame0 correction was implemented to recover stars saturated in the first group (21 s), but were unsaturated in the first 10.7 s comprising frame0 (ramp_fit.suppress_one_group=False). For Stage-2 output files (*_cal.fits), image1overf.py111https://github.com/chriswillott/jwst was used to correct for 1/f noise. Because the small FOV contained few suitable Gaia reference stars, astrometric alignment was done relatively between frames and filters. A simple source catalog was created from a single arbitrary exposure in the F115W filter using source detection from the JWST pipeline. Alignment of the remaining F115W frames and frames from the remaining filters to this source catalog was accomplished using the JWST/Hubble Alignment Tool (JHAT)222https://github.com/arminrest/jhat. Because alignment to Gaia was already accomplished via JHAT, the Stage-3 tweakreg step was skipped.
@@ 3.1 Data-reduction Methodology
JWST pipeline version 1.12.0 with CRDS version 11.17.9 and context jwst_1149.pmap were used in the creation of the final F770W imaging data. The remaining MIRI filters were subsequently processed using CRDS version 11.17.6 and context jwst_1179.pmap. The new CRDS context included updates for the MIRI Medium Resolution Spectrometer and thus does not affect the MIRI imaging data we present here. Flux-calibrated images in all four filters were produced by processing the raw MIRI ramp files using Dectector1Pipeline, with the output being run through Image2Pipeline with default parameters. Before combining the individual dithers into a final image, we aligned the F770W Image2Pipeline files (*_cal.fits) to the NIRCam F444W point-source catalog. We then applied the same astrometric corrections to the remaining MIRI filters. Following image alignment, we perform instrumental background subtraction following the method outlined in Nally et al. (2024). For each filter, we first median combine all dithers. Because I Zw 18 is present in each of these exposures, however, the median-combined backgrounds are not free of structure. We therefore select unaffected detector rows and columns and map these to the detector plane to produce a model background, which we then subtract from each individual exposure (see Nally et al., 2024; Dicken et al., 2024). Finally, the background-subtracted and F444W-aligned individual dither exposures were combined into final images using Image3Pipeline with the tweakreg step turned off.
3.2 Point-source Extraction
@@
The extraction of point sources in I Zw 18 was accomplished through use of the starbugii (Nally, 2023) photometric tool and pipeline. This software is optimized for JWST observations of crowded stellar fields within complex environments, performing point-source extraction and band merging across multiple observations and wavelengths, utilizing core functions from the python photutils (Bradley et al., 2022) package. Examples of programs which have employed starbugii for photometric study include Jones et al. (2023), Lenkić et al. (2024), Nally et al. (2024), and Habel et al. (2024). The complete set of the relevant starbugii parameters and their adopted values used for this program’s photometric extractions are listed in Table 3.
@@
For each filter in NIRCam, we perform source detection using starbug2 detect on the Stage-3 drizzled image. We locate sources that are above the sky level, and their location is measured by fitting centroids. Values of the sharpness and roundness are assigned en route to calculating a geometry for each point source. By setting an upper limit on sharpness, the measured ratio of source peak height to its median pixel value, remaining image artifacts are removed from the catalog, while setting a lower limit assists in removing faint peaks in the dust structure. Source symmetry is measured by the roundness parameter. Most resolved background galaxies, as well as further spurious detections within the dust structure, are removed by limiting the allowed level of asymmetry in the source profiles. The same routine was used to detect sources in the F770W MIRI image with modified detection parameters. Because the decrease in image resolution inhibited the detection of robust point-like sources in the longer-wavelength filters (F1000W, F1500W, and F1800W), bright sources in close proximity to other similar objects were instead merged into bright clustered regions that were not analyzed in this study.
@@
Aperture photometry was performed for all four NIRCam filters (F115W, F200W, F356W, and F444W), plus the shortest-wavelength MIRI filter (F770W). For all NIRCam sources, we use a fixed radius of 1.5 pixels as well as a sky annulus of 3.0 and 4.5 pixels for the inner and outer radii, respectively. Aperture correction is interpolated between values given in CRDS jwst_nircam_apcorr_0004.fits. For the MIRI F770W band, the procedure is the same, however the aperture radius is increased to 2 pixels to account for the larger PSFs intrinsic to longer wavelengths, backgrounds are calculated within annuli of 4 and 6 pixels for the inner and outer radii, respectively, and jwst_miri_apcorr_0005.fits is used to calculate aperture corrections. Through inspection of the data quality array within the aperture of each source, we flag those with saturated, DO_NOT_USE, or SRC_JMP pixels (indicating a jump during detection) with the SRC_BAD flag in starbugii.
@@
We perform PSF photometry on the stellar sources of I Zw 18 on the individual Stage-2 exposures of all four NIRCam filters. First, the nebulous emission which underlies our imaging data is modeled with the routine starbug2 background. This places masking apertures of variable size, and fills them with the median pixel value within a local annulus. Values of the diffuse emission background are then measured for every pixel in the image by averaging all local pixels within a set box size, creating an effective representation of the nebulous emission. A clean image of the field, with nebulous emission estimate subtracted, is therefore created which enables accurate fitting of PSFs.
The tool webbpsf (Perrin et al., 2014) version 1.1.1 was then employed to generate a 5 arcsec PSF for each NIRCam detector subarray. After estimating and subtracting the nebulous background from each single exposure, PSF photometry was performed by running starbug2 psf on the background-subtracted image, at locations defining source positions determined via the source detection. By leaving the centroid position as a free parameter, both the flux and position of each source are allowed to be fit. If the position of the new centroid differs from the initial guess determined during the source-detection step by more than 0″.1, the flux is refit with the source’s position held fixed to the initial guess and is flagged with SRC_FIX, denoting the fit as being of poorer quality.
The calculation and application of instrumental zero-point magnitudes are necessary to calibrate the PSF photometry, as they are not normalized to physical units. A median difference in source magnitudes measured by the aperture photometry and PSF fitting is calculated using starbug2 calc-instr-zp, allowing for the determination of instrumental zero-point magnitudes, which are then employed to calibrate the PSF photometry from starbugii to the AB magnitude system. These magnitudes are then converted to the Vega system using the reference file jwst_nircam_abvegaoffset_0002.asdf.
@@ 3.3 Stellar Source Catalog
Point-source catalogs for each of the four NIRCam filters were created via the methods described in §3.2. A band-matched catalog of point sources is then constructed for I Zw 18 by merging these catalogs produced for the individual filters using starbug2-match band. The possibility of mismatching high-quality stellar data with any spurious detector artifacts is mitigated by excluding all sources in the individual filter catalogs with an error on their magnitude greater than 10%. While this may limit the potential for completeness of the master catalog, this strategy allows for better reliability overall.
First, we consider the catalog made from the shortest-wavelength filter data, for which the smaller PSF FWHM possesses astrometry of the highest certainty. These sources are then matched using a nearest-neighbor algorithm to sources in the next-shortest-wavelength catalog. The threshold for separation between filter catalogs increases with increasing wavelength (equivalently, as the size of the PSF increases), as adopting a single separation threshold which is larger than the astrometric uncertainty of, for example, the longest-wavelength NIRCam filter would result in mismatches with, for example, the shortest-wavelength NIRCam filter. For NIRCam, the SW filters (F115W and F200W) are matched using a threshold of 0″.06, while the LW filters (F356W and F444W) are matched and added using a threshold of 0″.10. Any unmatched sources outside of the separation threshold are then appended to the end of the catalog, before the process is repeated with the next-shortest-wavelength filter data which remains. We note that, for the nearest-neighbor search method, the position coordinates for the source are adopted from the shortest-wavelength catalog from which said source initially appears. This allows us to retain the greatest possible positional accuracy for each source.
@@
To avoid artifacts at the edges of the detectors and unresolved background galaxies impacting our photometric analysis, an elliptical region approximately 55″40″ aligned along I Zw 18’s semi-major axis was defined. Only point sources detected within this region were considered to be members of I Zw 18 in our analysis. In addition, point sources from Component C were extracted from within a 30″22″ elliptical region centered on this system, while carefully avoiding the detector edge. The proximity of Component C to the detector edge makes robust separation between real sources and artifacts challenging; here, we present preliminary results in §4.3, with a more detailed analysis to follow as the JWST calibrations improve.